OpenStudy (anonymous):

Andrea can clean a house 4 times as fast as Denise. When they work together, Andrea and Denise can clean a large house in 8 hours. How many hours would it take Denise to clean the house by herself? 20 15 10 5

OpenStudy (anonymous):

15

OpenStudy (anonymous):

How do you know? @ariyanamckeown

OpenStudy (raden):

let x = Denise's time if do by herself 4x = Andrea's time if do by herself the time when they work together can be writtten as (4x * x)/(4x + x) = 8 4x^2/5x = 8 4x/5 = 8 4x = 5 * 8 4x = 40 x = 10 so, Denise's time id work by herself is 10 hours

OpenStudy (raden):

* if

OpenStudy (anonymous):

(A + D)8 = 1 A = 4D so, (4D + D)8 = 1 5D = 1/8 D = 1/40 so, it would take Denise 40 hours. The right answer is not on the list, unless you meant to ask for the hours it takes Andrea to clean the house